A 3.3
g sample of sodium hydrogen carbonate is added to a solution of acetic acid weighing 10.7
g. The two substances react, releasing carbon dioxide gas to the atmosphere. After the reaction, the contents of the reaction vessel weigh 11.9
g. What is the mass of carbon dioxide released during the reaction
1
Expert's answer
2019-09-18T05:29:51-0400

NaHCO3 + CH3COOH = CH3COONa + ^CO2 + H2O

3.3 + 10.7 = 14 g total mass.

14 - 11.9 = 2.1 g mass of CO2.